# Rydberg atomic polarimetry of radio-frequency fields

### v4

@@ I Introduction
Since the turn of the millennium, Rydberg atoms have experienced a veritable renaissance, emerging as a workhorse for quantum-technological applications [1]. Prompted by the proposal to exploit their long range dipolar interactions for quantum-computational gate operations [2], extensive efforts were devoted to making Rydberg atomic quantum information processors a reality[3]. Along another avenue, Rydberg atoms found themselves at the heart of vapor-based sensing applications [4, 5, 6, 7, 8] following the seminal 2012 work by Sedlacek et al.[9]
@@ II.1 Polarization dependent RF sensing in Rydberg EIT—a conceptual model
To set the scene for Rydberg atomic RF-polarimetry, we introduce in Fig. 1(a) the energy level diagram for a fictitious four-level model atom that includes angular momentum degeneracy for the ground and second excited levels. A probe laser field linearly polarized along the $z$-axis matches the optical transition frequency between the ground level $g$ and an excited level $i$; choosing $\bf\hat{z}$ as the quantization axis for the model atom, the probe field can drive a $\pi$-transition between $|g,m=0\rangle$ (one of three degenerate angular momentum states of the ground level) and $|i,m=0\rangle$.
@@
We next consider the effect of adding a linearly-polarized RF field to our model atom that resonantly couples the Rydberg level $r_{1}$ to another Rydberg level $r_{2}$. It is unimportant if $r_{2}$ lies above or below $r_{1}$ in energy and without loss of generality we will depict it above $r_{1}$ [see Fig. 1(a)]. For the RF field polarization aligned at an angle $\theta=0^{\circ}$ with respect to the optical fields, the non-degenerate state $|r_{2},m=0\rangle$ connects to $|r_{1},m=0\rangle$, the top rung of the EIT ladder. The RF coupling $|r_{2},m=0\rangle\!\!\leftrightarrow\!\!|r_{1},m=0\rangle$ acts to disrupt the optical EIT for $\Delta_{\text{c}}=0$ through the Autler-Townes (AT) effect [29]. For $\theta=90^{\circ}$, the $|r_{1},m=0\rangle$ top rung of the EIT ladder is left unconnected by the RF-field and EIT proceeds exactly as if the RF field had not been present. Figure 1c shows level diagrams and the coupled states for the two cases $\theta=0^{\circ},90^{\circ}$; the state-splitting resulting from the coupling is indicated with dashed lines. Figure 1d displays the associated EIT spectra of the optical probe transmission versus the coupling laser field detuning $\Delta_{\text{c}}$. This is double-peaked (AT doublet) for $\theta=0^{\circ}$ and single-peaked for $\theta=90^{\circ}$ while a triple-peaked spectrum is found at $\theta=45^{\circ}$ (all spectra for angles between $\theta=0^{\circ}$ and $\theta=90^{\circ}$ will have three peaks, with the height ratio between the central peak and a side lobe increasing with $\theta$).
Figure 2: Hyperfine level diagrams for the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{5/2}\!\!\leftrightarrow\!\!P_{3/2}$ excitation ladder with the ultimate and penultimate levels (Rydberg levels) resonantly coupled by an RF field, which is linearly polarized along the atomic quantization axis and drives $\pi$-transitions (orange arrows). Red and blue arrows shows the allowed transitions for optical probe and coupling fields linear polarized (a) perpendicularly to the RF field (optical $\sigma$-transitions) and (b) parallel to the RF field (optical $\pi$-transitions). The $F$-levels for the Rydberg states are energetically degenerate, and have been offset vertically to illustrate the dipole-allowed transitions in play.
@@ II.2 Rydberg atomic RF-polarimetry in a complex multi-level atom
Figure 3: Hyperfine level diagrams for the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{3/2}\!\!\leftrightarrow\!\!P_{1/2}$ excitation ladder with the ultimate and penultimate levels (Rydberg levels) resonantly coupled (orange arrows) by an RF field, which is linearly polarized along the atomic quantization axis. Optical coupling (blue arrows) and probe (red arrows) fields are also linearly polarized along the quantization axis and $\pi$-transitions are driven throughout the system.
@@
Figure 4: RF-dressed Rydberg levels—procedure for finding the energy structure probed by EIT. (a) A linearly-polarized RF field couples states of lower ($J=5/2$, black) and upper ($J=3/2$, grey) Rydberg levels via $\pi$-transitions with dipole matrix elements as indicated. (b) Field-dressed Rydberg levels with $m_{J}$-dependent AT-splitting. Each grey-black state in the diagram represents an equal two-component superposition of $J=5/2$ and $J=3/2$ states with the adjacent $\pm$ indicating its symmetry. (c) Labelling the lower dressed manifold of (b) to include all possible values for the nuclear spin projection $m_{I}$ so that each state inside the hexagon is designated by $|m_{J},m_{I}\rangle$ (and hence $m_{F}=m_{J}+m_{I}$) as well as a symmetry parameter $s$. Each $|m_{J},m_{I};s\rangle$ can be expanded using (2) which facilitates the evaluation of the transition strength for the optical coupling from a state $|J^{\prime}=3/2,F^{\prime}=3,m_{F}^{\prime}=m_{F}\rangle$ through (3).
@@
The scenario of parallel RF and optical polarizations is captured in Fig. 2(b)[32]. Here Ref. 24 argues that “In this case, $\pi$ transitions are driven throughout the system and all excitation pathways experience a four level system. The theoretical and experimental spectra have two transmission peaks…”. Our theoretical treatment will, as we shall see, in fact predict four transmission peaks for a $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{5/2}\!\!\leftrightarrow\!\!P_{3/2}$ ladder system; the reason that only two peaks are observed in Ref. 24 is because insufficient RF power is applied to resolve the structure in the presence of line broadening mechanisms. But more importantly and at the heart of understanding the physics in play: the absence of a central transmission peak is not the result of $\pi$-transitions being driven throughout the system! As a case in point, we will contrast the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{5/2}\!\!\leftrightarrow\!\!P_{3/2}$ system (which we shall call type-I) to that of $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{3/2}\!\!\leftrightarrow\!\!P_{1/2}$ (type-II). Figure 3 shows a level diagram of the latter with transitions for co-polarized RF and optical fields. Similarly to the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{5/2}\!\!\leftrightarrow\!\!P_{3/2}$ system in Fig. 2b, only $\pi$-transitions are driven throughout the system, yet, as we shall demonstrate, the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{3/2}\!\!\leftrightarrow\!\!P_{1/2}$ system displays a dominant central transmission peak for $\theta=0^{\circ}$. In fact, the central transmission peak encountered at $\theta=0^{\circ}$ dwarfs the central peak at $\theta=90^{\circ}$.
@@ II.3.1 Type-I system: absent central EIT peak for co-polarized fields
To seek actual the origin of the absent central EIT peak of the system of Fig. 2(b)
@@
Figure 5: Diagram for evaluating the optical coupling from a state $|J^{\prime}=3/2,F^{\prime}=3,m_{F}^{\prime}=m_{F}\rangle$ to a state in a field-dressed $D_{3/2}\!\!\leftrightarrow\!\!P_{1/2}$ Rydberg manifold. The procedure for obtaining this diagram follows that of Fig. 4 and each of the states $|m_{J},m_{I};s\rangle$ can be expanded using (2).
@@
To establish how the RF field dressing of the atom affects the optical EIT probing we transform to a hyperfine basis, where $J$ is coupled to the spin of the atomic nucleus $I=3/2$ to form $F=J+I$. Figure 4(c) recasts the lower dressed manifold of Fig. 4(b) (orange hexagon) to include the possible projections $m_{I}$ of the nuclear spin. In doing so, each state of Fig. 4(c) is designated by quantum numbers $m_{J}$ and $m_{I}$ in addition to a parameter $s=0,\pm 1$ that tracks the symmetry of the parental dressed state of Figure 4(b) from which it descends: every dressed state in Fig. 4(b) is either an equal symmetric ($s=+1$) or antisymmetric (s=1s=-1) superposition of $r_{1}$ and $r_{2}$ components. The $r_{1}$ spectator states that are unaffected by the dressing field ($m_{J}=\pm 5/2$) are labeled by $s=0$. Expanding the $r_{1}$ part (i.e., $J=5/2$) of each state of Fig. 4(c) as
@@
we obtain the required transformation to the hyperfine basis. A measure for the strength of an optical $\pi$-transition from a particular state $|J^{\prime}F^{\prime}m_{F}^{\prime}\rangle$ of the $i$-level (i.e., $J^{\prime}=3/2$) to a state of the lower dressed Rydberg manifold can now be found (see Supplementary Note 1):
@@
The first 3-$j$ symbol of (3) vanishes if $|m_{J}|>J^{\prime}$. An optical field co-polarized with the linear RF dressing field can therefore not couple any of states of the $P_{J^{\prime}=3/2}$ $i$-level manifold to the dressed Rydberg states with $m_{J}=\pm 5/2$. The EIT spectrum will hence not display a central peak.
@@ II.3.2 Type-II system: dominant central EIT peak for co-polarized fields
presented in Fig. 3 yields the dressed level diagram shown in Fig. 5. When, for this type-4 ladder, we consider the dipole matrix elements for optical $\pi$-transitions from the $i$-level to the Rydberg manifold, we see that $|m_{J}|=\{\tfrac{1}{2},\tfrac{3}{2}\}\leq J^{\prime}=\tfrac{3}{2}$. We do therefore not experience the vanishing of first 33-$j$ symbol in (3). Rather, its common, non-zero squared magnitude
@@ III Experiment
Figure 6(c) shows the transmitted probe light for a frequency scan of the coupling light over the two transitions we utilize for RF-sensing (see appendix A). Both probe and coupling beams are linearly polarized in the horizontal plane. A photomixing antenna (see appendix B) emits linearly-polarized $\sim 55$ GHz RF-radiation down onto the vapour cell along a vertical axis intersecting the horizontally propagating optical beams. At the point of intersection, the angle between polarizations of the RF field and the optical fields is $\theta$ (see Figure 6a). We can vary this angle by rotating the photoconductive antenna about its axis.
Figure 7: (a) Spectrograms from EIT-probing the RF-dressed $34D_{5/2}\!\!\leftrightarrow\!\!35P_{3/2}$ (left column) and $35D_{3/2}\!\!\leftrightarrow\!\!36P_{1/2}$ (right column) transitions. The transmission of the probe field is measured on a parameter space spanned by the RF-polarization angle $\theta$ and the coupling laser detuning $\Delta_{\text{c}}$ (cf. Fig. 6). From the top row, which is a reference spectrogram with no applied RF-field, moving down the RF-power increases; the RF-power shows an monotonous albeit nonlinear dependence with the bias voltage $V_{\text{Bias}}$ of the photomixing antenna. (b) Transmitted probe light on the $\Delta_{\text{c}}=0$ axis of the $V_{\text{Bias}}=-1.5$ V spectrograms as a function $\theta$ with as a function of $\theta_{\text{MW}}$ (blue lines); sinusoidal fits shown as dashed orange lines.
Figure 6(d) shows EIT spectra acquired for the resonantly driven $34D_{5/2}\!\!\leftrightarrow\!\!35P_{3/2}$ (left panel, type-1 ladder) and $35D_{3/2}\!\!\leftrightarrow\!\!36P_{1/2}$ (right panel, type-4 ladder) Rydberg transitions, for $\theta=0^{\circ}$, $45^{\circ}$, and $90^{\circ}$. The polarization signatures for the type-1 ladder in the left panel are similar to those reported for the $53D_{5/2}\!\!\leftrightarrow\!\!54P_{3/2}$ Rydberg transition in Ref. 24. In particular, for $\theta=0^{\circ}$ we note the absence of a central peak and what appears to be an AT doublet. Our dressed-picture analysis of the type-1 system in Sec. II.3.1 predicted a five peak spectrum with no central component (an effective four peak spectrum) at $\theta=0^{\circ}$. The differential splitting of the Doppler broadened $m_{J}=\pm 3/2$ and $m_{J}=\pm 5/2$ lines is however too small at our applied RF power to be resolved. Rotating the emitter to $\theta=90^{\circ}$, we observe, again in correspondence with Ref. 24, how the heights of the split peaks reduce while a prominent central peak emerges at $\Delta_{\text{c}}=0$. Also for this situation, each of the split peaks in fact contains two spectral features that are obscured by broadening.
Figure 8: Simulated spectrograms (bottom row) for type-I (left column) and type-II ladders (right column). The corresponding experimental measurements ($V_{\text{Bias}}=-0.5$ V) are shown for comparision (top row).

### v5

# EIT Spectroscopy of Atoms with RF-Dressed Rydberg Levels: Role of Quantized Angular Momentum

@@ I Introduction
Since the turn of the millennium, Rydberg atoms have experienced a veritable renaissance, emerging as a workhorse for quantum-technological applications [1]. Prompted by the proposal to exploit their long range dipolar interactions for quantum-computational gate operations [2], extensive efforts were devoted to making Rydberg atomic quantum information processors a reality [3]. Along another avenue, Rydberg atoms found themselves at the heart of vapor-based sensing applications [4, 5, 6, 7, 8] following the seminal 2012 work by Sedlacek et al. [9].
@@ II.1 Polarization dependent RF sensing in Rydberg EIT—a conceptual model
To elucidate how an EIT signal might depend on the polarization of an applied RF field, we introduce in Fig. 1(a) the energy level diagram for a fictitious four-level model atom that includes angular momentum degeneracy for the ground and second excited levels. A probe laser field linearly polarized along the $z$-axis matches the optical transition frequency between the ground level $g$ and an excited level $i$; choosing $\bf\hat{z}$ as the quantization axis for the model atom, the probe field can drive a $\pi$-transition between $|g,m=0\rangle$ (one of three degenerate angular momentum states of the ground level) and $|i,m=0\rangle$.
@@
We next consider the effect of adding a linearly-polarized RF field to our model atom that resonantly couples the Rydberg level $r_{1}$ to another Rydberg level $r_{2}$. It is unimportant if $r_{2}$ lies above or below $r_{1}$ in energy and without loss of generality we will depict it above $r_{1}$ [see Fig. 1(a)]. For the RF field polarization aligned at an angle $\theta=0^{\circ}$ with respect to the optical fields, the non-degenerate state $|r_{2},m=0\rangle$ connects to $|r_{1},m=0\rangle$, the top rung of the EIT ladder. The RF coupling $|r_{2},m=0\rangle\!\!\leftrightarrow\!\!|r_{1},m=0\rangle$ acts to disrupt the optical EIT for $\Delta_{\text{c}}=0$ through the Autler-Townes (AT) effect [29]. For $\theta=90^{\circ}$, the $|r_{1},m=0\rangle$ top rung of the EIT ladder is left unconnected by the RF-field and EIT proceeds exactly as if the RF field had not been present. Figure 1(c) shows level diagrams and the coupled states for the two cases $\theta=0^{\circ},90^{\circ}$; the state-splitting resulting from the coupling is indicated with dashed lines. Figure 1(d) displays the associated EIT spectra of the optical probe transmission versus the coupling laser field detuning $\Delta_{\text{c}}$. This is double-peaked (AT doublet) for $\theta=0^{\circ}$ and single-peaked for $\theta=90^{\circ}$ while a triple-peaked spectrum is found at $\theta=45^{\circ}$ (all spectra for angles between $\theta=0^{\circ}$ and $\theta=90^{\circ}$ will have three peaks, with the height ratio between the central peak and a side lobe increasing with $\theta$).
Figure 2: Hyperfine level diagrams for the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{5/2}\!\!\leftrightarrow\!\!P_{3/2}$ excitation ladder with the ultimate and penultimate levels (Rydberg levels) resonantly coupled by an RF field, which is linearly polarized along the atomic quantization axis and drives $\pi$-transitions (orange arrows). Red and blue arrows show the allowed transitions for optical probe and coupling fields linear polarized (a) perpendicularly to the RF field (optical $\sigma$-transitions) and (b) parallel to the RF field (optical $\pi$-transitions). The $F$-levels for the Rydberg states are energetically degenerate, and have been offset vertically to illustrate the dipole-allowed transitions in play.
@@ II.2 Rydberg atomic RF-polarimetry in a complex multi-level atom
Figure 3: Hyperfine level diagrams for the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{3/2}\!\!\leftrightarrow\!\!P_{1/2}$ excitation ladder with the ultimate and penultimate levels (Rydberg levels) resonantly coupled (orange arrows) by an RF field, which is linearly polarized along the atomic quantization axis. The optical coupling (blue arrows) and probe (red arrows) fields are also linearly polarized along the quantization axis and $\pi$-transitions are driven throughout the system.
@@
Figure 4: RF-dressed Rydberg levels—procedure for finding the energy structure probed via EIT. (a) A linearly-polarized RF field couples states of lower ($J=5/2$, black) and upper ($J=3/2$, grey) Rydberg levels via $\pi$-transitions with dipole matrix elements as indicated. (b) Field-dressed Rydberg levels with $m_{J}$-dependent AT-splitting. Each grey-black state in the diagram represents an equal two-component superposition of $J=5/2$ and $J=3/2$ states with the adjacent $\pm$ indicating its symmetry. (c) Labelling the lower dressed manifold of (b) to include all possible values for the nuclear spin projection $m_{I}$ so that each state inside the hexagon is designated by $|m_{J},m_{I}\rangle$ (and hence $m_{F}=m_{J}+m_{I}$) as well as a symmetry parameter $s$. Each $|m_{J},m_{I};s\rangle$ can be expanded using (2) which facilitates the evaluation of the transition strength for the optical coupling from a state $|J^{\prime}=3/2,F^{\prime}=3,m_{F}^{\prime}=m_{F}\rangle$ through (3).
@@
The scenario of parallel RF and optical polarizations is captured in Fig. 2(b) [32]. Here, Ref. 24 argues that “In this case, $\pi$ transitions are driven throughout the system and all excitation pathways experience a four level system. The theoretical and experimental spectra have two transmission peaks…”. Our theoretical treatment will, as we shall see, in fact predict four transmission peaks for a $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{5/2}\!\!\leftrightarrow\!\!P_{3/2}$ ladder system; the reason that only two peaks are observed in Ref. 24 is because insufficient RF power is applied to resolve the structure in the presence of line broadening mechanisms. But more importantly and at the heart of understanding the physics in play: the absence of a central transmission peak is not the result of $\pi$-transitions being driven throughout the system! As a case in point, we will contrast the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{5/2}\!\!\leftrightarrow\!\!P_{3/2}$ system (which we shall call type-I) to that of $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{3/2}\!\!\leftrightarrow\!\!P_{1/2}$ (type-II). Figure 3 shows a level diagram of the latter with transitions for co-polarized RF and optical fields. Similarly to the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{5/2}\!\!\leftrightarrow\!\!P_{3/2}$ system in Fig. 2b, only $\pi$-transitions are driven throughout the system, yet, as we shall demonstrate, the $S_{1/2}^{F=2}\!\!\leftrightarrow\!\!P_{3/2}^{F=3}\!\!\leftrightarrow\!\!D_{3/2}\!\!\leftrightarrow\!\!P_{1/2}$ system displays a dominant central transmission peak for $\theta=0^{\circ}$. In fact, the central transmission peak encountered at $\theta=0^{\circ}$ dwarfs the central peak at $\theta=90^{\circ}$.
@@ II.3.1 Type-I system: absent central EIT peak for co-polarized fields
To seek the actual, underlying origin of the absent central EIT peak of the system of Fig. 2(b)
@@
Figure 5: Diagram for evaluating the optical coupling from a state $|J^{\prime}=3/2,F^{\prime}=3,m_{F}^{\prime}=m_{F}\rangle$ to a state in a field-dressed $D_{3/2}\!\!\leftrightarrow\!\!P_{1/2}$ Rydberg manifold. The procedure for obtaining this diagram follows that of Fig. 4 and each of the states $|m_{J},m_{I};s\rangle$ can be expanded using Eq. (2).
@@
To establish how the RF field dressing of the atom affects the optical EIT probing, we transform to a hyperfine basis, where $J$ is coupled to the spin of the atomic nucleus $I=3/2$ to form $F=J+I$. Figure 4(c) recasts the lower dressed manifold of Fig. 4(b) (orange hexagon) to include the possible projections $m_{I}$ of the nuclear spin. In doing so, each state of Fig. 4(c) is designated by quantum numbers $m_{J}$ and $m_{I}$ in addition to a parameter $s=0,\pm 1$ that tracks the symmetry of the parental dressed state of Figure 4(b) from which it descends: every dressed state in Fig. 4(b) is either an equal symmetric ($s=+1$) or antisymmetric (s=1s=-1) superposition of $r_{1}$ and $r_{2}$ components. The $r_{1}$ spectator states that are unaffected by the dressing field ($m_{J}=\pm 5/2$) are labeled by $s=0$. Expanding the $r_{1}$ part (i.e., $J=5/2$) of each state of Fig. 4(c) as
@@
we obtain the required transformation to the hyperfine basis.
Using the expansion 2, a measure for the strength of an optical $\pi$-transition from a particular state $|J^{\prime}F^{\prime}m_{F}^{\prime}\rangle$ of the $i$-level (i.e., $J^{\prime}=3/2$) to a state of the lower dressed Rydberg manifold can now be found (see Supplementary Note 1):
@@
The first 3-$j$ symbol of Eq. (3) vanishes if $|m_{J}|>J^{\prime}$. An optical field co-polarized with the linear RF dressing field can therefore not couple any of states of the $P_{J^{\prime}=3/2}$ $i$-level manifold to the dressed Rydberg states with $m_{J}=\pm 5/2$—the central states in Fig. 4(c). As as result, the EIT spectrum obtained from probing the RF-dressed Rydberg states will not display a central peak.
@@ II.3.2 Type-II system: dominant central EIT peak for co-polarized fields
presented in Fig. 3 yields the dressed level diagram shown in Fig. 5. When, for this type-4 ladder, we consider the dipole matrix elements for optical $\pi$-transitions from the $i$-level to the Rydberg manifold, we see that $|m_{J}|=\{\tfrac{1}{2},\tfrac{3}{2}\}\leq J^{\prime}=\tfrac{3}{2}$. We do therefore not experience the vanishing of first 33-$j$ symbol in Eq. (3). Rather, its common, non-zero squared magnitude
@@ III Experiment
Figure 6(c) shows the transmitted probe light for a frequency scan of the coupling light over the $i\leftrightarrow r_{1}$ transitions we utilize in our EIT sensing scheme (see appendix A). Both probe and coupling beams are linearly polarized in the horizontal plane. A photomixing antenna (see appendix B) emits linearly-polarized $\sim 55$ GHz RF-radiation down onto the vapour cell along a vertical axis intersecting the horizontally propagating optical beams. At the point of intersection, the angle between polarizations of the RF field and the optical fields is $\theta$ (see Figure 6a). We can vary this angle by rotating the photoconductive antenna about its axis.
Figure 7: (a) Spectrograms from EIT-probing the RF-dressed $34D_{5/2}\!\!\leftrightarrow\!\!35P_{3/2}$ (left column) and $35D_{3/2}\!\!\leftrightarrow\!\!36P_{1/2}$ (right column) transitions. The transmission of the probe field is measured on a parameter space spanned by the RF polarization angle $\theta$ and the coupling laser detuning $\Delta_{\text{c}}$ (cf. Fig. 6). From the top row, which is a reference spectrogram with no applied RF-field, moving down the RF power increases; the RF power shows a monotonous albeit nonlinear dependence with the bias voltage $V_{\text{Bias}}$ of the photomixing antenna. (b) Transmitted probe light on the $\Delta_{\text{c}}=0$ axis of the $V_{\text{Bias}}=-1.5$ V spectrograms as a function $\theta$ with as a function of $\theta_{\text{MW}}$ (blue lines); sinusoidal fits shown as dashed orange lines.
Figure 6(d) shows EIT spectra acquired for the resonantly driven $34D_{5/2}\!\!\leftrightarrow\!\!35P_{3/2}$ (left panel, type-1 ladder) and $35D_{3/2}\!\!\leftrightarrow\!\!36P_{1/2}$ (right panel, type-4 ladder) Rydberg transitions, for $\theta=0^{\circ}$, $45^{\circ}$, and $90^{\circ}$. The polarization signatures for the type-1 ladder in the left panel are similar to those reported for the $53D_{5/2}\!\!\leftrightarrow\!\!54P_{3/2}$ Rydberg transition in Ref. 24. In particular, for $\theta=0^{\circ}$ we note the absence of a central peak and what appears to be an AT doublet. Our dressed-picture analysis of the type-1 system in Sec. II.3.1 predicted effective four peak spectrum (a five peak spectrum with no central component) at $\theta=0^{\circ}$. The differential splitting of the Doppler broadened $m_{J}=\pm 3/2$ and $m_{J}=\pm 5/2$ lines is however too small at our applied RF power to be resolved. Rotating the emitter to $\theta=90^{\circ}$, we observe, again in correspondence with Ref. 24, how the heights of the split peaks reduce while a prominent central peak emerges at $\Delta_{\text{c}}=0$. Also for this situation, each of the split peaks in fact contains two spectral features that are obscured in the experiments by broadening.
Figure 8: Simulated spectrograms (bottom row) for type-I (left column) and type-II ladders (right column). The corresponding experimental measurements ($V_{\text{Bias}}=-1.5$ V) are shown for comparision (top row).